# Quasiprobability Thermodynamic Uncertainty Relation

### v2

@@ Abstract
where $\rho$ is the density operator, $H$ the Hamiltonian, $L_{k}$ jump operators, $\cdot$$\cdot$] the commutator, and $\{\cdot,\cdot\}$ the anticommutator. We refer to $\mathcal{L}$ and $\mathcal{D}$ as the Liouvillian and the dissipator, respectively. While we assume the Hamiltonian and jump operators are time-independent, our results are easily extended to time-dependent cases. Given the initial data $\rho_{0}$, Eq. (2) is formally solved by $\rho(t)=e^{\mathcal{L}t}\rho_{0}$. We indicate the adjoint of $\mathcal{L}$ and $\mathcal{D}$ by $\mathcal{L}^{\dagger}$ and $\mathcal{D}^{\dagger}$. For an observable (Hermitian operator) $X$, we describe the eigendecomposition as $X=\sum_{x}x\Pi_{x}$ with eigenvalues $x$ and projectors $\Pi_{x}$. Its expectation value at time $t$ is given by $\langle X\rangle_{t}=\mathop{\mathrm{tr}}\nolimits(X\rho(t))=\sum_{x}xp(x,t)$, where $p(x,t)=\mathop{\mathrm{tr}}\nolimits(\Pi_{x}\rho(t))$. We divide the time derivative into the Hamiltonian part $J_{X}^{\mathrm{H}}(t)=i\langle[H,X]\rangle_{t}$ and the dissipative part $J_{X}^{\mathrm{d}}(t)=\langle\mathcal{D}^{\dagger}(X)\rangle_{t}$ as $d_{t}\langle X\rangle_{t}=J_{X}^{\mathrm{H}}(t)+J_{X}^{\mathrm{d}}(t)$.
@@
Theoretical studies have revealed that the moment generating function can be directly measured by devising an interferometric scheme [43, 42], which has been experimentally realized for $X=H$ [66]. Thus, $\langle(\Delta X)^{n}\rangle_{t,t+\Delta t}$ is not only theoretically but also practically a reasonable measure of the dynamical fluctuation in a quantum object $X$. Additionally, as shown in End Matter, $\mathsf{G}_{t}(\lambda,\Delta t)$ can be obtained by appropriately “quantizing” the classical counterpart.
@@
Combining it with Eq. (11), we obtain the TUR (10).
@@
Note that the negativity of the quasiprobability also indicates the genuine quantumness called contextuality [46, 47]. The other one (Q2) is also non-classical because it means the chance of escape grows more rapidly than the number of evacuation targets increases. The statement shows that either of these quantumness conditions has to hold in any basis for the anomalous scaling to occur.
When $\rho$ commutes with $X$, the quasiprobability conditions necessitate non-classicality at the level of basis components for every simultaneous eigenbasis. Now, we define an eigenbasis $\{|s,j\rangle\}$ to be classical if the following conditions are satisfied for all $(s,j)$ and $(s^{\prime},j^{\prime})(\neq(s,j))$: as $N$ increases, (1) the jump rate does not grow ($|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|=O(1)$), and (2) the number of jumps that mediate the states does not grow ($\{k\colon|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|>0\}=O(1)$). Then, if a simultaneous eigenbasis $\{|s,j\rangle\}$ of $\rho$ and $X$ is classical, we can prove that the anomalous scaling does not occur. First observe that $R_{sjsj}=\sum_{s^{\prime},j^{\prime}(\neq s,j)}\sum_{k}|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|^{2}=O(N)$, and thus $\bar{\lambda}(\rho)=-\sum_{s,j}R_{sjsj}p_{sj}=O(N)$, where $p_{s,j}=\mathop{\mathrm{tr}}\nolimits(\rho\Pi_{s,j})$. Besides, we have $T_{s^{\prime}j^{\prime}sj}(\rho)=\sum_{k}|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|^{2}p_{s,j}\geq 0$ since $\rho$ is diagonal with respect to $\{|s,j\rangle\}$. Thus, $m_{X}(\rho)$ can grow at most at $O(N)$ because both (Q1) and (Q2) are violated by the eigenbasis. This is in accordance with the result of Ref. [57], where dissipationless current is proved not to occur when $\rho$ has no coherence with respect to this type of eigenbasis.
@@
Let us prove the main claim that either (Q1) or (Q2) is necessary for anomalous scaling. First, because $\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)=\bar{\lambda}(\rho)$, the negation of (Q2) implies that $\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)$ is $O(N)$. Then, split the summation as $\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)=\mathcal{T}^{+}-\mathcal{T}^{-}$ with $\mathcal{T}^{+}=\sum_{s,s^{\prime}|\mathcal{T}_{s^{\prime}s}(\rho)\geq 0}\mathcal{T}_{s^{\prime}s}(\rho)$ and $\mathcal{T}^{-}=\sum_{s,s^{\prime}|\mathcal{T}_{s^{\prime}s}(\rho)<0}|\mathcal{T}_{s^{\prime}s}(\rho)|$. Due to the negation of (Q1), the negative part $\mathcal{T}^{-}$ is at most $O(N)$. Therefore, the positive part $\mathcal{T}^{+}$ can also be at most $O(N)$. Finally, $m_{X}(\rho)$ is bounded as
$m_{X}(\rho)$ $\leq\sum_{s,s^{\prime}}(x_{s}-x_{s^{\prime}})^{2}|\mathcal{T}_{s^{\prime}s}(\rho)|$
$\leq\max_{s,s^{\prime}|\mathcal{T}_{s^{\prime}s}(\rho)\neq 0}(x_{s}-x_{s^{\prime}})^{2}(\mathcal{T}^{+}+\mathcal{T}^{-}),$
and thus is of order $N$.
@@ Acknowledgements.
Acknowledgments.— The authors are grateful to Hiroyasu Tajima and Ken Funo for fruitful discussions. K.Y. is supported by the Special Postdoctoral Researchers Program at RIKEN and JSPS KAKENHI Grant No. 24H00834. R.H. is supported by JSPS KAKENHI Grant No. JP24K16982. K.Y. and R.H. are supported by JST ERATO Grant No. JPMJER2302, Japan.
@@
Appendix A Derivation of Eq. (11)
@@
Appendix B Computational details of Example
@@
Appendix C Hierarchy of conditions

### v3

@@ Abstract
where $\rho$ is the density operator, $H$ the Hamiltonian, $L_{k}$ jump operators, $\cdot$$\cdot$] the commutator, and $\{\cdot,\cdot\}$ the anticommutator. We refer to $\mathcal{L}$ and $\mathcal{D}$ as the Liouvillian and the dissipator, respectively. While we assume the Hamiltonian and jump operators are time-independent, our results are easily extended to time-dependent cases. Given the initial data $\rho_{0}$, Eq. (2) is formally solved by $\rho(t)=e^{\mathcal{L}t}\rho_{0}$. We indicate the adjoint of $\mathcal{L}$ and $\mathcal{D}$ by $\mathcal{L}^{\dagger}$ and $\mathcal{D}^{\dagger}$. For an observable (Hermitian operator) $X$, we describe the eigendecomposition as $X=\sum_{x}x\Pi_{x}$ with eigenvalues $x$ and projectors $\Pi_{x}$. While we focus on time-independent observables for simplicity, our results are valid for time-dependent ones (detailed later). Its expectation value at time $t$ is given by $\langle X\rangle_{t}=\mathop{\mathrm{tr}}\nolimits(X\rho(t))=\sum_{x}xp(x,t)$, where $p(x,t)=\mathop{\mathrm{tr}}\nolimits(\Pi_{x}\rho(t))$. We divide the time derivative into the Hamiltonian part $J_{X}^{\mathrm{H}}(t)=i\langle[H,X]\rangle_{t}$ and the dissipative part $J_{X}^{\mathrm{d}}(t)=\langle\mathcal{D}^{\dagger}(X)\rangle_{t}$ as $d_{t}\langle X\rangle_{t}=J_{X}^{\mathrm{H}}(t)+J_{X}^{\mathrm{d}}(t)$.
@@
Theoretical studies have revealed that the moment generating function can be directly measured by devising an interferometric scheme [43, 42], which has been experimentally realized for $X=H$ [66]. Thus, $\langle(\Delta X)^{n}\rangle_{t,t+\Delta t}$ is not only theoretically but also practically a reasonable measure of the dynamical fluctuation in a quantum object $X$. Additionally, as shown in Supplemental Material [67], $\mathsf{G}_{t}(\lambda,\Delta t)$ can be obtained by appropriately “quantizing” the classical counterpart.
@@
In End Matter, we prove that Eq. (8) is valid for $n=2$ even if $X$ depends on time.
@@
Combining it with Eq. (11), we obtain the TUR (10). Note that this discussion applies to time-dependent observables since Eq. (8) remains valid.
@@
Note that the negativity of the quasiprobability also indicates the genuine quantumness called contextuality [46, 47]. The other one (Q2) is also non-classical because it means the chance of escape grows more rapidly than the number of evacuation targets increases  [^2] In other words, if we adopt conditions (1) $O(1)$ transition rates and (2) $O(1)$ channel multiplicity, which are detailed two paragraphs later, as the definition of classicality, the escape rate scales at most $O(N)$ in classical situations.. The statement shows that either of these quantumness conditions has to hold in any basis for the anomalous scaling to occur.
The necessity of either (Q1) or (Q2) for anomalous scaling is proved by showing that it is equivalent to the existence of a pair $(s,s^{\prime})$ such that $|\mathcal{T}_{s^{\prime}s}(\rho)|/N\to\infty$ as $N\to\infty$ (for the proof, see End Matter). Under this equivalence, it becomes evident that if the conditions (Q1) and (Q2) are violated, we do not observe anomalous scaling since $m_{X}(\rho)\leq\sum_{s,s^{\prime}}(x_{s}-x_{s^{\prime}})^{2}|\mathcal{T}_{s^{\prime}s}|$. We stress that the split of the formal condition into (Q1) and (Q2) enables us to bridge anomalous scaling and the nonclassicality of quasiprobabilities.
When $\rho$ commutes with $X$, the quasiprobability conditions necessitate non-classicality at the level of basis components for every simultaneous eigenbasis. Now, we define an eigenbasis $\{|s,j\rangle\}$ to be classical if the following conditions are satisfied for all $(s,j)$ and $(s^{\prime},j^{\prime})(\neq(s,j))$: as $N$ increases, (1) the jump rate does not grow ($|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|=O(1)$), and (2) the number of jumps that mediate the states does not grow ($\{k\colon|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|>0\}=O(1)$)  [^3] These conditions on bare transition rates will be naturally satisfied if we consider few-body systems or systems with only short-range interactions.. Then, if a simultaneous eigenbasis $\{|s,j\rangle\}$ of $\rho$ and $X$ is classical, we can prove that the anomalous scaling does not occur. First observe that $R_{sjsj}=\sum_{s^{\prime},j^{\prime}(\neq s,j)}\sum_{k}|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|^{2}=O(N)$, and thus $\bar{\lambda}(\rho)=-\sum_{s,j}R_{sjsj}p_{sj}=O(N)$, where $p_{s,j}=\mathop{\mathrm{tr}}\nolimits(\rho\Pi_{s,j})$. Besides, we have $T_{s^{\prime}j^{\prime}sj}(\rho)=\sum_{k}|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|^{2}p_{s,j}\geq 0$ since $\rho$ is diagonal with respect to $\{|s,j\rangle\}$. Thus, $m_{X}(\rho)$ can grow at most at $O(N)$ because both (Q1) and (Q2) are violated by the eigenbasis. This is in accordance with the result of Ref. [57], where dissipationless current is proved not to occur when $\rho$ has no coherence with respect to this type of eigenbasis.
@@ Acknowledgements.
Acknowledgments.— The authors are grateful to Hiroyasu Tajima, Ken Funo, and Yuuya Chiba for fruitful discussions. K.Y. is supported by the Special Postdoctoral Researchers Program at RIKEN and JSPS KAKENHI Grant No. 24H00834. R.H. is supported by JSPS KAKENHI Grant No. JP24K16982. K.Y. and R.H. are supported by JST ERATO Grant No. JPMJER2302, Japan.
@@
Appendix A Quantization of generating function
We show that the moment generating function (6) is obtained by appropriately quantizing the classical counterpart. Consider classical Markov processes with discrete mesostates $i=1,2,\dots,n$. The system’s state is described by a probability density $\bm{p}=(p_{i})\in\mathbb{R}_{\geq 0}^{n}$ such that $p_{i}\geq 0$ and $\sum_{i}p_{i}=1$. Its time evolution is generally written as the classical master equation
$\frac{d\bm{p}(t)}{dt}=\mathcal{L}_{\mathrm{cl}}\big(\bm{p}(t)\big)\coloneqq R\bm{p}(t),$ (S1)
where $R$ is the rate matrix, which we assume time-independent [65]. Given the initial probability density $\bm{p}_{0}$, the classical master equation (S1) is solved by $\bm{p}(t)=e^{Rt}\bm{p}_{0}=e^{\mathcal{L}_{\mathrm{cl}}t}\bm{p}_{0}$.
Let us consider a quantity that depends on the system’s mesostate, $\bm{f}=(f_{i})$. Its expectation value at time $t$ is provided as $\langle\bm{f}\rangle_{t}=\sum_{i}f_{i}p_{i}(t)$. By using the standard inner product $\langle\bm{v},\bm{w}\rangle\coloneqq\sum_{i}v_{i}w_{i}$, it is rewritten as $\langle\bm{f}\rangle_{t}=\langle\bm{f},\bm{p}(t)\rangle$. The Heisenberg picture can be discussed by defining $\bm{f}(t)=e^{R^{\mathrm{T}}t}\bm{f}$, where $\mathrm{T}$ indicates transpose. Because $(e^{Rt})^{\mathrm{T}}=e^{R^{\mathrm{T}}t}$, we find $\langle\bm{f}(t)\rangle_{0}=\langle\bm{f}\rangle_{t}$. The adjoint of $\mathcal{L}_{\mathrm{cl}}$ with respect to $\langle\cdot,\cdot\rangle$ is given by $\mathcal{L}_{\mathrm{cl}}^{\dagger}(\bm{f})=R^{\mathrm{T}}\bm{f}$. It determines the time evolution of $\bm{f}(t)$ as $\frac{d}{dt}\bm{f}(t)=\mathcal{L}_{\mathrm{cl}}^{\dagger}(\bm{f}(t))$.
We next consider the statistics of the increment of $\bm{f}$. The joint probability that the system is in state $i$ at time $t$ and in $j$ at time $t+\Delta t$ is given by $e^{R\Delta t}]_{ji}p_{i}(t) [$65]. Therefore, the change in $\bm{f}$, denoted by $\Delta\bm{f}$, has the moments
$\big\langle(\Delta\bm{f})^{n}\big\rangle_{t,t+\Delta t}=\sum_{i,j}(f_{j}-f_{i})^{n}[e^{R\Delta t}]_{ji}p_{i}(t).$ (S2)
As shown later, they are generated by the moment generating function defined as
$\mathsf{G}_{t}^{\mathrm{cl}}(\lambda,\Delta t)\coloneq\Big\langle e^{\mathcal{L}_{\mathrm{cl}}^{\dagger}\Delta t}(e^{i\lambda\bm{f}})e^{-i\lambda\bm{f}}\Big\rangle_{t},$ (S3)
where the exponential and product of vectors are interpreted entrywise; $e^{\bm{v}}=(e^{v_{i}})$ and $\bm{v}\bm{w}=(v_{i}w_{i})$ for any vectors $\bm{v}$ and $\bm{w}$. That is, we can prove
$\big\langle(\Delta\bm{f})^{n}\big\rangle_{t,t+\Delta t}=(-i\partial_{\lambda})^{n}\mathsf{G}_{t}^{\mathrm{cl}}(\lambda,\Delta t)|_{\lambda=0}.$ (S4)
Importantly, the TMH moment generating function (6) is led to by the following replacement
$\mathcal{L}_{\mathrm{cl}}\to\mathcal{L},\;\;\bm{f}\to X,\;\;\bm{p}\to\rho,$ (S5)
and the introduction of the anticommutator between $e^{\mathcal{L}^{\dagger}\Delta t}(e^{i\lambda X})$ and $e^{-i\lambda X}$. If the anticommutator is not applied, $\mathsf{G}_{t}^{\mathrm{cl}}$ will become the moment generating function corresponding to the Kirkwood–Dirac quasiprobability [42].
Let us prove Eq. (S4). First, expand $\mathsf{G}_{t}^{\mathrm{cl}}$ as
$\mathsf{G}_{t}^{\mathrm{cl}}(\lambda,\Delta t)$
$=\sum_{m=0}^{\infty}\sum_{l=0}^{\infty}\frac{(i\lambda)^{m}}{m!}\frac{(i\lambda)^{l}}{l!}\big\langle e^{\mathcal{L}^{\dagger}\Delta t}(\bm{f}^{m})(-\bm{f})^{l}\big\rangle_{t}$
$=\sum_{r=0}^{\infty}\sum_{l=0}^{r}\frac{(i\lambda)^{r}}{r!}\binom{r}{l}\big\langle e^{\mathcal{L}^{\dagger}\Delta t}(\bm{f}^{r-l})(-\bm{f})^{l}\big\rangle_{t}.$
Differentiating by $\lambda$ for $n$ times and making $\lambda$ zero, we finally find the terms corresponding to $r=n$ in the summation; so we get
$(-i\partial_{\lambda})^{n}\mathsf{G}_{t}^{\mathrm{cl}}(\lambda,\Delta t)|_{\lambda=0}$
$=\sum_{l=0}^{n}\binom{n}{l}\big\langle e^{\mathcal{L}^{\dagger}\Delta t}(\bm{f}^{n-l})(-\bm{f})^{l}\big\rangle_{t}$
$=\sum_{l=0}^{n}\binom{n}{l}\sum_{i,j}[e^{R\Delta t}]_{ji}f_{j}^{n-l}(-f_{i})^{l}p_{i}(t)$
$=\sum_{i,j}[e^{R\Delta t}]_{ji}p_{i}(t)\sum_{l=0}^{n}\binom{n}{l}f_{j}^{n-l}(-f_{i})^{l}$
$=\sum_{i,j}(f_{j}-f_{i})^{n}[e^{R\Delta t}]_{ji}p_{i}(t)=\langle(\Delta\bm{f})^{n}\rangle_{t,t+\Delta t},$
which concludes the proof.
Let us focus on the second moment in the short-time limit. By expanding $e^{R\Delta t}$, we find
$\big\langle(\Delta\bm{f})^{2}\big\rangle_{t,t+\Delta t}=\sum_{i,j}(f_{j}-f_{i})^{2}R_{ji}p_{i}(t)\Delta t+O(\Delta t^{2}),$ (S6)
where the zeroth order term vanishes because $(f_{j}-f_{i})^{n}\delta_{ji}p_{i}=0$ for any $i$ and $j$. Thus, the short-time second moment $m_{\bm{f}}^{\mathrm{cl}}(t)=\lim_{\Delta t\to 0}\big\langle(\Delta\bm{f})^{2}\big\rangle_{t,t+\Delta t}/\Delta t$ reads
$m_{\bm{f}}^{\mathrm{cl}}(t)$ $=\sum_{i,j}(f_{j}-f_{i})^{2}R_{ji}p_{i}(t)$ (S7)
$=\sum_{i,j}f_{j}^{2}R_{ji}p_{i}(t)-2\sum_{i,j}f_{i}f_{j}R_{ji}p_{i}(t),$ (S8)
where we used $\sum_{j}R_{ji}=0$. By remembering $\mathcal{L}_{\mathrm{cl}}^{\dagger}=R^{\mathrm{T}}$, we can further rewrite this equation into
$m_{\bm{f}}^{\mathrm{cl}}(t)=\big\langle\mathcal{L}_{\mathrm{cl}}^{\dagger}(\bm{f}^{2})-2\mathcal{L}_{\mathrm{cl}}^{\dagger}(\bm{f})\bm{f}\big\rangle_{t}.$ (S9)
We can quantize this equation and recover Eq. (13) by $\mathcal{L}^{\dagger}(\bm{f})\bm{f}\to\{\mathcal{L}^{\dagger}(X),X\}/2$ in addition to the replacement in Eq. (S5).
Appendix B Derivation of Eq. (11)
@@ Appendix A Quantization of generating function
Appendix C Computational details of Example
@@
Appendix D Hierarchy of conditions